OpenStudy (anonymous):

What are the x– and y–intercepts of y = 2x2 – 10x + 8? answer please

OpenStudy (anonymous):

do u only need answers or explaination as well?

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

is tht 2x^2?

OpenStudy (anonymous):

y-intercept is (0, 8)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x^2 – 10x + 8 = 0 x^2 - 5x + 4 = 0 (x - 4) (x - 1) = 0 x = 4 x = 1 are the x-intercepts